At a concession stand, 3 hot dogs and 4 hamburgers cost $11.00; 4 hot dogs and 3 hamburgers cost $10.00. find the cost of one hot dog and the cost of one hamburger.
Answer by TimothyLamb(4379)

x = cost of dog
y = cost of burger
---
3x + 4y = 11
4x + 3y = 10
---
put the system of linear equations into standard form
---
3x + 4y = 11
4x + 3y = 10
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = cost of dog = $1
y = cost of burger = $2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php